# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 5.2.4 Direct and Converse Opt-Lyapunov Theorems
The function ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and continuous. Thus, we have, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(V(T^{k}(x))$. By Def. 9 and Prop. 17, $V(T^{k}(x))\leq(N_{V}(T))^{k}V(x)\leq N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing, we conclude that $\nu_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}\right)$. Note that $x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$ is strictly increasing and continuous on $J\supset[0,1]$ as ${{\alpha}_{\left(J\right)}^{-1}}$ is and $\overline{V_{X^{\mathrm{in}}}}$ is not zero. Moreover, by Prop. 18 and hypothesis, $N_{V}(T)\in(0,1)$. As $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, there exists $j\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{j})>0$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and $\alpha(I)\ni 0$, $\nu_{j}>\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(0)=\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(0\times\overline{V_{X^{\mathrm{in}}}}\right)=h(0)$. ∎
Example 10.
Following Th. 9, from the Opt-Lyapunov function $V$ defined at Example 5 and the ($X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, $\alpha$ defined at Example 9, the function $h:[0,1]\ni s\mapsto\alpha_{\mathbb{R}}^{-1}(s\overline{V_{X^{\mathrm{in}}}})$ belongs to $\mathbb{\Gamma}_{\rm fun}(\nu)$ and $\overline{N_{T}}(V)=4/9\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$. Then $h_{V}(s)=s+p^{2}/3-\min\{\zeta,\varepsilon\}$ with $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$ and $\zeta\in(0,p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})$ where $n_{1/2}^{*}$ is the maximizer for the function $n\mapsto\varphi_{n}(1/2)$.
We draw a table as we done at Example 3 to compare the direct (from the function $h$ of Example 3) and the Opt-Lyapunov approaches. We have some flexibility about $\zeta$ so we propose two values for a comparison purpose. Hence, we compare the computations for $h_{\zeta_{1}}$ associated to $\zeta_{1}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/2$ and $h_{\zeta_{2}}$ associated to $\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$. The evaluations for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},4/9)\rfloor$ and for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},4/9)\rfloor$ are given at Table 3.
$p=30$ $p=9$ $p=3$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mathbf{K}^{s}_{\nu}$ 8 11 22 5 8 19 2 3 17
$\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ 9 12 23 7 9 20 3 7 18
$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},\frac{4}{9})\rfloor$ 8 14 36 5 11 33 4 8 31
$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},\frac{4}{9})\rfloor$ 8 14 36 9 11 33 12 12 31
Table 3: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},\beta)\rfloor$ described at Eq. (4) for the sequence $\nu$ of Eq. (2) and two pairs $(h_{\zeta_{1}},4/9)$ and $(h_{\zeta_{2}},4/9)$ of Example 10.
The evaluations of $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},4/9)\rfloor$ from the Opt-Lyapunov function proposed at Example 5 and associated $(X^{\mathrm{in}},T,\varphi)$-certificates of compatibility provide worst upper bounds when $\mu$ is small (equal to $1/10$ or $1/1000$) but the optimal value for $\mu=1/3$. A small value for $\zeta$ ($\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$) makes decrease the precision for $\mu=1/3$. We recall that the y-intercept of the affine function $h_{\zeta}$ is $p^{2}/3-\min\{\varepsilon,\eta\}$ where $\varepsilon$ is a very small value when $\mu$ is small. Hence, when $\mu=1/10$ or $\mu=1/1000$, $\zeta_{1}$ or $\zeta_{2}$ are still greater than $\varepsilon$.
Note that the values $h_{\zeta_{i}}^{-1}(\nu_{k})$ are strictly positive if and only if for $k\in\operatorname{Argmax}(\nu)$. This is sufficient to compute an upper bound of $\mathbf{K}^{s}_{\nu}$. ∎
We establish a ”converse Lyapunov theorem”. The construction of the Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is inspired by the one proposed by Yoshizawa [Yos66].
@@
Let $(h,\beta)\in\mathbb{\Gamma}(\nu)$ be useful for $\nu$. Let us define $s\mapsto\omega(s):=\min\left\{h(1),\max\left\{s,h(0)\right\}\right\}$. Then, the function $V\in\mathcal{PSD}$ defined as follows:
$x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega\left(\varphi(T^{k}(x))\right)\right)$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Moreover, the function $\hat{h}$ defined as follows:
@@
Let us prove that $V$ takes its values in $[0,+\infty]$. Let $x\in\mathbb{R}^{d}$. As we have $h(1)\geq\omega(s)\geq h(0)$, we have, as ${{h}_{\left([0,1]\right)}^{-1}}$ is increasing, $1\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\omega(\varphi(T^{k}(x))))\geq 0$ for all $k\in\mathbb{N}$. We conclude, as $\beta\in(0,1)$ that $V(x)\geq 0$ and $V(x)=+\infty$ for all $x\in\mathbb{R}^{d}$ such that $\varphi(T^{k}(x))\geq h(1)$ for all $k$ sufficiently large.
By definition of $\mathbb{\Gamma}(\nu)$, we have for all $x\in X^{\mathrm{in}}$ and all $n\in\mathbb{N}$, $\varphi(T^{n}(x))\leq h(\beta^{n})<h(1)$ and so $V(x)\leq\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(\beta^{k}))=\sup_{k\in\mathbb{N}}\beta^{-k}\beta^{k}=1$. This implies that $\overline{V_{X^{\mathrm{in}}}}\in[0,1]$. By assumption, $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and $\mathcal{S}(\nu,h)$ is nonempty. Let us take $k\in\mathbb{N}$ such that $h(1)>\nu_{k}>h(0)$. So there exists $\varepsilon>0$ such that $\nu_{k}-\varepsilon>h(0)$. For this $\varepsilon$, there exists $x\in X^{\mathrm{in}}$ such that $\nu_{k}\geq\varphi(T^{k}(x))\geq\nu_{k}-\varepsilon$. Finally, for this $x$, $V(x)\geq\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\varphi(T^{k}(x))>\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$ as ${{h}_{\left([0,1]\right)}^{-1}}$ strictly increases. We conclude that $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$. Finally, we have to prove that $V\circ T\leq\beta V$. Let $y\in\mathbb{R}^{d}$, if $V(y)=+\infty$, the result holds. Now suppose that $y\in\operatorname{dom}V$, we have:
$\begin{array}[]{rl}V(T(y))=&\displaystyle{\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{k+1}(y))\right)}\\ =&\displaystyle{\sup_{j\in\mathbb{N}^{*}}\beta^{-j+1}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ =&\displaystyle{\beta\sup_{j\in\mathbb{N}^{*}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ \leq&\displaystyle{\beta\sup_{j\in\mathbb{N}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)=\beta V(y)}\end{array}$
@@ 6 A Technical Summary
We have proposed a function $\mathfrak{F}_{\nu}$ at Eq. (4) whose the integer part of values provide solutions for Problem 2. This function is based on couples $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$. Without this condition, $(h,\beta)$ is useless as $\mathfrak{F}_{\nu}$ is equal to +$\infty$+ ∞. When the supremum of $(\nu_{k})_{k}$ is strictly greater than its limit superior, a useful couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ exists (Th. 1) and $h$ can be chosen as an affine function. Furthermore, we can compute the greatest maximizer of $(\nu_{k})_{k}$ from this couple (Th. 3) and thus we can solve Problem 1. Unfortunately, to directly compute a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is as difficult as solving Problem 1 and we have to construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from alternative methods. We propose alternatives coming from classical tools from stability theory of difference equations.
@@
The inequality $\varphi(T^{k}(x))\leq h(\beta^{k})$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ can be viewed as a kind of $\mathcal{KL}$ inequality that appears in asymptotic stability studies. Indeed, $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$ making the sequence $(h(\beta^{k}))_{k}$ strictly decreasing and convergent to $h(0)$. From this observation, we propose a more general $\mathcal{K}\mathcal{L}$ class of functions called $\mathcal{KL}_{\rm gen}$ where the positivity with respect to the state variable and the continuity assumptions are removed (compared with the $\mathcal{K}\mathcal{L}$ class). We have proved (Th. 5) that a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ exists if and only if $\nu_{\rm opt}$ is finite. Then, we extend Sontag’s result to the $\mathcal{KL}_{\rm gen}$ class (Lemma 2 and Lemma 3) to construct $h\in\Omega([0,1])$, useful for $\nu$, from a $\mathcal{KL}_{\rm gen}$ function and such that $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ (Th. 6). One of the advantage of the Sontag’s result extension is that the scalar part of elements of $\mathbb{\Gamma}(\nu)$ can be fixed and it remains to construct $h$ from the $\mathcal{KL}_{\rm gen}$ certificate. In practice, to decide whether a function $(\gamma,\operatorname{Fin}\left(X^{\mathrm{in}}\right))$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$ seems difficult to make.
@@
One constructive alternative of $\mathcal{K}\mathcal{L}$ certificates consists in Lyapunov functions. Classical Lyapunov functions satisfy pure functional inequations and are not defined from the orbits of the dynamical system. One of the most interesting theoretical result for Lyapunov functions is the necessity of their existence for the stability (namely converse Lyapunov theorems). For our specific problem, we propose a new type of Lyapunov function called Opt-Lyapunov functions for which we obtain both direct (Th. 9) and converse theorem (Th 10). The main difference between classical and Opt-Lyapunov functions is that Opt-Lyapunov functions are nonnegative but not definite (many zeros) and can take infinite values (outside reachable values taken by the dynamical system). The fact that Opt-Lyapunov functions are not definite (and not continuous) makes classical comparison functions impossible to use. In the classical case, comparison functions serve as link between our objective function $\varphi$ and the classical Lyapunov function. This link (and comparison functions) is replaced by the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion (Th. 8). Theorem 11 sums up the equivalences proved in this paper relatively to the peaks computation problem.
@@
The following statements are equivalent:
@@
    There exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)$ is finite and $\nu_{\rm opt}=\max_{k\in\mathbb{N}}\{\nu_{k}:k=0,\ldots,\mathbf{K}^{s}_{\nu}\}$;
@@
The proof of Theorem is done by sketching the functions constructed in the proofs of earlier theorems. Figure 1 provides a graphical view of different results involved to prove Th. 11. In Figure 1, the notation $\mathcal{OL}_{\rm comp}$ stands for the set of $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov functions.
@@
As $V\in\mathcal{OL}_{\rm comp}$ is $(X^{\mathrm{in}},T,\varphi)$-compatible, we associate with them a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. For a such a function $V$, we denote by $\alpha_{V}$ the chosen $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
We now clarify the circled numbers that appear in Figure 1.
@@
    The function $h$ cannot be explicitly formulated and is constructed following the proofs of Lemma 2 and Lemma 3.
@@
    where $h_{V}$ is defined as follows:
    $[0,1]\ni s\mapsto h_{V}(s)=\displaystyle{{\alpha_{V}}_{\left(J\right)}^{-1}}(s\overline{V_{X^{\mathrm{in}}}})$
@@
      $(-\infty,h(1)]\ni s\mapsto\alpha_{V}(s)=\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in[h(0),h(1)]\end{array}\right.$
Appendix
Proof of Lemma 4.
Lemma 4.

### v4

@@ 5.3 Direct and Converse Opt-Lyapunov Theorems
The function ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and continuous on $[0,1]$. Thus, we have, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$,
$\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(V(T^{k}(x)).$
By Def. 5 and Prop. 5, $V(T^{k}(x))\leq(N_{V}(T))^{k}V(x)\leq N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing, we conclude that
$\nu_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}\right).$
Note that $x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$ is strictly increasing and continuous on $J\supset[0,1]$ as ${{\alpha}_{\left(J\right)}^{-1}}$ is, and $\overline{V_{X^{\mathrm{in}}}}$ is not zero. Moreover, by Prop. 6 and hypothesis, $N_{V}(T)\in(0,1)$. As $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, there exists $j\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{j})>0$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and $\alpha(I)\ni 0$, $\nu_{j}>\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(0)=\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(0\times\overline{V_{X^{\mathrm{in}}}}\right)=h(0)$. ∎
We establish a converse Lyapunov theorem. The construction of the Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is inspired by the one proposed by Yoshizawa [Yos66].
@@
Let $(h,\beta)\in\mathbb{\Gamma}(\nu)$ be useful for $\nu$. Let us define
$\mathbb{R}\ni s\mapsto\mathfrak{m}(s):=\min\left\{h(1),\max\left\{s,h(0)\right\}\right\}.$
Then, the function $V\in\mathcal{PSD}$ is defined as follows:
$\mathbb{R}^{d}\ni x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}\left(\varphi(T^{k}(x))\right)\right)$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Moreover, the function $\hat{h}$ is defined as follows:
@@
Let us prove that $V$ takes its values in $0,+\infty]$. Let $x\in\mathbb{R}^{d}$. As we have for all $s\in\mathbb{R}$ $h(1)\geq\mathfrak{m}(s)\geq h(0)$, we have, as ${{h}_{\left([0,1]\right)}^{-1}}$ is increasing, $1\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\mathfrak{m}(\varphi(T^{k}(x))))\geq 0$ for all $k\in\mathbb{N}$. We conclude, as $\beta\in(0,1)$ that $V(x)\geq 0$ and $V(x)=+\infty$ for all $x\in\mathbb{R}^{d}$ such that $\varphi(T^{k}(x))\geq h(1)$ for all $k$ sufficiently large.
By definition of $\mathbb{\Gamma}(\nu)$, we have for all $x\in X^{\mathrm{in}}$ and all $n\in\mathbb{N}$, $\varphi(T^{n}(x))\leq h(\beta^{n})<h(1)$ and so
$V(x)\leq\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(\beta^{k}))=\sup_{k\in\mathbb{N}}\beta^{-k}\beta^{k}=1.$
This implies that $\overline{V_{X^{\mathrm{in}}}}\in[0,1]$. By assumption, $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and $\mathcal{S}(\nu,h)$ is nonempty. Let us take $k\in\mathbb{N}$ such that $h(1)>\nu_{k}>h(0)$. Therefore, there exists $\varepsilon>0$ such that $\nu_{k}-\varepsilon>h(0)$. For this $\varepsilon$, there exists $x\in X^{\mathrm{in}}$ such that $\nu_{k}\geq\varphi(T^{k}(x))\geq\nu_{k}-\varepsilon$. Finally, for this $x$,
$V(x)\geq\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\varphi(T^{k}(x))>\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$
as ${{h}_{\left([0,1]\right)}^{-1}}$ strictly increases. We conclude that $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$. Finally, we must prove that $V\circ T\leq\beta V$. Let $y\in\mathbb{R}^{d}$. If $V(y)=+\infty$, there is nothing to prove. Assume now $V(y)<+\infty$. We have:
$\begin{array}[]{rl}V(T(y))=&\displaystyle{\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{k+1}(y)))\right)}\\ =&\displaystyle{\sup_{j\in\mathbb{N}^{*}}\beta^{-j+1}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{j}(y)))\right)}\\ =&\displaystyle{\beta\sup_{j\in\mathbb{N}^{*}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{j}(y)))\right)}\\ \leq&\displaystyle{\beta\sup_{j\in\mathbb{N}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\mathfrak{m}(\varphi(T^{j}(y)))\right)=\beta V(y)}\end{array}$
@@
$\begin{array}[]{l}V(T^{k}(x))=\\ \sup_{n\in\mathbb{N}}\beta^{-n}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\max\left\{\varphi(T^{k+n}(x)),h(0)\right\}\right)\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\max\left\{\varphi(T^{k}(x)),h(0)\right\}\right)\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\varphi(T^{k}(x))\right)\end{array}$
if $\varphi(T^{k}(x))\geq h(0)$. If $\varphi(T^{k}(x))<h(0)$, then $V(T^{k}(x))\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$ and $\hat{h}(\varphi(T^{k}(x)))\leq 0$. We conclude that $\hat{h}(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$. As $\mathcal{S}(\nu,h)\neq\emptyset$, there exists $k\in\mathbb{N}$, such that $h(1)>\nu_{k}>h(0)$. As ${{h}_{\left([0,1]\right)}^{-1}}$ is strictly increasing on $[h(0),h(1)]$, we have $\hat{h}(\nu_{k})=\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\nu_{k})>0$. This ends the proof that $\hat{h}$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility of $V$. ∎
Example 10.
As we choose to take $\gamma$ equals
$\max\left\{\dfrac{2(\omega-\eta)^{-3}}{(M(1+\omega^{-2})},\dfrac{V(x^{*})}{\eta}\right\},$
the function
$\alpha:s\mapsto\left\{\begin{array}[]{lr}\dfrac{C}{M{}(1+s^{-2})}&\text{ if }s\geq z\\ \gamma(s-z)+\dfrac{C}{M^{2}(1+z^{-2})}&\text{ if }s<z\end{array}\right$.
is a $(X_{\ell},L,\varphi_{\ell})$-certificate of compatibility for $V_{\ell}$. Finally, $(h,\beta)\in\mathbb{\Gamma}(\omega)$ where:
$h=\alpha^{-1}\text{ and }\beta=\dfrac{1}{\min\{a^{2}+c^{2},b^{2}\}}.$
This leads to the formula for all $k\in\mathbb{N}$ such that $\omega_{k}>\alpha^{-1}(0)$
$\dfrac{\ln(\alpha(\omega_{k})/\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}})}{-\ln(\min\{a^{2}+c^{2},b^{2}\})}.$
Applying our computations to the numerical data of Subsection 3.4 leads to $\alpha^{-1}(0)=\dfrac{1757}{64}>8=\omega_{0}$. Then, we cannot use $\alpha$ at $\omega_{0}$. Finally,
$\dfrac{\ln(\alpha(\omega_{1})/\overline{{V_{\ell}}_{X^{\mathrm{in}}_{\ell}}})}{-\ln(\min\{a^{2}+c^{2},b^{2}\})}=\dfrac{\ln\left(\dfrac{49}{22592}\right)}{\ln\left(\dfrac{4}{5}\right)}=27.49$
This means that from the certificate of compatibility, we assume that the maximal term of $\omega$ is one of the 28th first terms. This result is sound but not optimal. Indeed, the maximal term is $\omega_{1}$.
We see that the Opt-Lyapunov approach is less precise than $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ approach. For the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ approach, we have to construct an upper bound of $\varphi_{\ell}(L^{k}x)$. This bas been done directly from the function $H$ that characterizes $\varphi_{\ell}(L^{k}x)$. This seems to explain why the upper bound is tight and thus the function $h\in\mathbb{\Gamma}(\omega)$ obtained provides better results. Opt-Lyapunov function $V_{\ell}$ defined in Eq. (32) does not depend on $\varphi_{\ell}$. The function depends only on the dynamics $L$ and initial conditions set. The link between $V_{\ell}$ and $\varphi_{\ell}$ is established in Example 8 and becomes explicit in Example 9. More generally, a certificate of compatibility is the function that relates a Lyapunov function to the orbits $\varphi(T^{k}x))$. ∎
@@ 6 A Technical Summary
In this paper, we have proposed a general method to solve optimization problems of the form (DPCP). More precisely, focusing on the integer component of the optimal solution, it is necessary to compute an index $k\in\mathbb{N}$ for which the term of the sequence $\nu$ (defined in (1)) is maximal. The method is based on a function $\mathfrak{F}_{\nu}$ defined in [Adj25c] and quickly recalled in Eq. (4). This function $\mathfrak{F}_{\nu}$ requires a pair $(h,\beta)$ where $h$ is a strictly increasing continuous function on $[0,1]$ and $\beta\in(0,1)$ such that $\nu_{k}\leq h(\beta^{k})$ for all $k\in\mathbb{N}$. A key notion is the usefulness of $(h,\beta)$, that is, there exists some $k\in\mathbb{N}$ such that $\nu_{k}>h(0)$. Without this condition, an upper bound of the greatest index maximizer of $\nu$ cannot be computed from $(h,\beta)$. However, the existence theorem [Adj25c, Th.2] (partially recalled in Th 1) is not totally constructive and computing $(h,\beta)\in\mathbb{\Gamma}(\nu)$ has to follow an alternative methodology. We propose alternatives based on classical tools from the stability theory of difference equations. In [Adj25a], we compute $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from classical $\mathcal{K}\mathcal{L}$ and Lyapunov functions. However, this approach suffers from a lack of converse theorems. Indeed, some problems of the form (DPCP) (such as Problem ($DPCP_{\ell}$)) cannot be solved using these classes of functions.
@@
The inequality $\varphi(T^{k}(x))\leq h(\beta^{k})$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ can be viewed as a kind of $\mathcal{KL}$ inequality that appears in asymptotic stability studies. Indeed, $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$ making the sequence $(h(\beta^{k}))_{k}$ strictly decreasing and convergent to $h(0)$. From this observation, we propose a more general $\mathcal{K}\mathcal{L}$ class of functions called $\mathcal{KL}_{\rm gen}$ where the positivity with respect to the state variable and the continuity assumptions are removed (compared with the $\mathcal{K}\mathcal{L}$ class). We have proved (Th.  2) that a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ exists if and only if $\nu_{\rm opt}$ is finite: Then, we extend Sontag’s result to the $\mathcal{KL}_{\rm gen}$ class (Lemma 1 and Lemma 2) to construct $h\in\Omega([0,1])$, useful for $\nu$, from a $\mathcal{KL}_{\rm gen}$ function and such that $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ (Th. 3). One of the advantages of the extension of Sontag’s result is that the scalar part of the elements of $\mathbb{\Gamma}(\nu)$ can be fixed, and it remains to construct $h$ from the $\mathcal{KL}_{\rm gen}$ certificate. In practice, to decide whether a function $(\gamma,\operatorname{Fin}\left(X^{\mathrm{in}}\right))$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$ seems difficult to make.
@@
One constructive alternative to $\mathcal{K}\mathcal{L}$ certificates consists in Lyapunov functions. Classical Lyapunov functions satisfy pure functional inequations and are not defined from the orbits of the dynamical system. One of the most interesting theoretical results for Lyapunov functions is the necessity of their existence for stability (namely, the converse Lyapunov theorems). For our specific problem, we propose a new type of Lyapunov function called the Opt-Lyapunov function, for which we obtain both direct (Th.  5) and the converse theorem (Th 6). The main difference between classical and Opt-Lyapunov functions is that Opt-Lyapunov functions are nonnegative but not definite (many zeros) and can take infinite values (outside reachable values taken by the dynamical system). The fact that Opt-Lyapunov functions are not definite (and not continuous) renders classical comparison functions impossible to use. In the classical case, comparison functions serve as a link between our objective function $\varphi$ and the classical Lyapunov function. This link (and comparison functions) is replaced by the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion (Th. 4). Theorem 7 summarizes the equivalences proved in this study with respect to the peak computation problem.
@@
The following statements are equivalent.
@@
    There exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)$ is well-defined and finite and $\nu_{\rm opt}=\max_{k\in\mathbb{N}}\{\nu_{k}:k=0,\ldots,\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)\}$;
@@
The proof of Theorem is obtained by sketching the functions constructed in the proofs of the earlier theorems. Figure 2 provides a graphical view of the different results involved in proving Th.  7. In Figure 2, the notation $\mathcal{OL}_{\rm comp}$ denotes the set of $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov functions.
@@
As $V\in\mathcal{OL}_{\rm comp}$ is $(X^{\mathrm{in}},T,\varphi)$-compatible, we associate with them a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. For such a function $V$, we denote by $\alpha_{V}$ the chosen $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
We now clarify the circled numbers in Figure 2.
@@
    The function $h$ cannot be explicitly formulated and is constructed following the proofs of Lemmas  1 and  2.
@@
    where $h_{V}$ is defined as
    $0,1]\ni s\mapsto h_{V}(s)=\displaystyle{{\alpha_{V}}_{\left(J\right)}^{-1}}(s\overline{V_{X^{\mathrm{in}}}})$
@@
[…]
Appendix A
A.1 Additional proofs for Opt-Lyapunov definition
We provide details about the proofs of Lemmas  3 and  4.
Lemma A.3.